OpenStudy (anonymous):

Will metal and fan! If x-3y=-9 and 5x-2y=7, what is the value of x? Options: A.1 B.2 C.3 D.4 Information-I am in Algebra 1(high school) and I have tried this so many times. I cannot find the correct answer!!!! Please help me!

OpenStudy (anonymous):

multiply the 5(x-3y=-9) so you can get rid of x by using elimination

OpenStudy (anonymous):

-17y=-52

OpenStudy (anonymous):

that should be correct. im not using a calculator so my math might be off.

OpenStudy (anonymous):

@iGreen. ??? Do you know this one? I am still confused

OpenStudy (anonymous):

Add 3y to both sides in the first equation. x - 3y = -9 x = 3y - 9 Now plug in 3y - 9 for x in the 2nd equation: 5(3y - 9) - 2y = 7. Now can you solve for y?